# Hyperparameter Optimization in Machine Learning

### v1

@@ Abstract
Hyperparameters are configuration variables controlling the behavior of machine learning algorithms. They are ubiquitous in machine learning and artificial intelligence and the choice of their values determine the effectiveness of systems based on these technologies. Manual hyperparameter search is often unsatisfactory and becomes unfeasible when the number of hyperparameters is large. Automating the search is an important step towards automating machine learning, freeing researchers and practitioners alike from the burden of finding a good set of hyperparameters by trial and error. In this survey, we present a unified treatment of hyperparameter optimization, providing the reader with examples and insights into the state-of-the-art. We cover the main families of techniques to automate hyperparameter search, often referred to as hyperparameter optimization or tuning, including random and quasi-random search, bandit-, model- and gradient-based approaches. We further discuss extensions, including online, constrained, and multi-objective formulations, touch upon connections with other fields such as meta-learning and neural architecture search, and conclude with open questions and future research directions.
@@
Over the last decade, machine learning has established itself as a key driver of innovation in science and industry. The digitalization of human interactions has led to the generation of data at an unprecedented scale. These data could be stored ever more cheaply thanks to advances in micro-electronics, as well as the advent of cloud computing. Progress in hardware and low-power chip design also led to an exponential increase of computing capabilities of respectively cloud infrastructure, such as graphical processing units (GPUs), and devices on the edge, such as mobile phones. These two trends, cheaper and better hardware for storage and compute, lay the foundation for the successful application of statistical machine learning in practice and its adoption in artificial intelligence. It enabled learning algorithms to sift through vast amounts of data to discover and extract patterns that can then be used to make data-driven decisions without or with limited human intervention.
@@
Perhaps the most notable illustration of the recent advances made in machine learning is AlphaGo [silver:2017:mastering_go], a computer program developed by machine learning experts at the London-based company DeepMind. AlphaGo won against one of the world champions of the game of Go, Lee Sedol, in 2016. This event generated considerable press coverage (and was even made into a movie) as it was believed at the time that no computer program could win a game of Go against a human before several decades to come. What is less known is that the success of AlphaGo critically depended on how a set of configuration parameters, called hyperparameters, were adjusted automatically by another computer program [chen:2018:bayesian_optimization_alphago]. This computer program relied on Bayesian optimization, an algorithm that sequentially predicts and then evaluates the performance of algorithms such as AlphaGo when it uses specific hyperparameters. Until a certain cost or time budget is exhausted, Bayesian optimization repeatedly computes an exploration-exploitation trade-off to determine what is the next most interesting set of hyperparameters to evaluate based on its performance observed so far.
@@
In practice, the generalization capabilities of any machine learning algorithm also depends on hyperparameters. Whereas conventional machine learning algorithms, such as support vector machines [shawe-taylor:2004:kernel_methods_pattern_analysis], need hand-engineered features to transform the raw data in a suitable format, deep neural networks learn in an “end-to-end” fashion by ingesting the raw data directly [lecun2015nature]. However, in both cases, the composition of the model (e.g., the choice of feature extractors or the neural network architecture) and the optimization algorithm depend on quantities that will determine how effective the algorithm can learn a specific natural language understanding, computer vision or speech recognition task.
@@
To illustrate the practical importance of carefully selecting the hyperparameters, we consider the problem of sentiment analysis. yogotama2015-emnlp15 studied the effect of hyperparameters in this context. More specifically, the authors framed the problem as a standard binary classification problem as is commonly done in the literature, where the task of the classifier is to predict whether the text expresses a negative or a positive sentiment. They compared simple logistic regression trained with stochastic gradient descent to convolutional neural networks, which achieved state-of-the-art results at the time of the publication. Table 1.1(a) shows the hyperparameters that were searched over. These included the choice of text features used (e.g., removal of stop words or not), the type of regularization (e.g., $\ell_{1}$ or $\ell_{2}$) and optimization algorithm parameters (e.g., convergence tolerance). The experimental results they reported on the Amazon electronics dataset are reproduced in Table 1.1(b). Interestingly, they showed that tuned logistic regression performed on par with a convolutional neural network using a bag of words representation of the text and only slightly worse that a sequential convolutional neural network. We will come back to this example in more details in the next chapter.
@@
Figure 1.1: We reproduce the results obtained by yogotama2015-emnlp15. Table 1.1(a) are the hyperparameters that were optimized. The feature extraction depends on the range of $n$-grams considered through $n_{min}$ and $n_{max}$. Stop words could be removed or not and words weighted according different weighting schemes. Two choices of regularization were considered and the amount of regularization was optimized. The convergence tolerance was also optimized. The binary classifiers compared are either SVM (support vector machine), LR (logistic regression), NN (multi-layer perceptron) or CNN (convolutional neural network). The authors found that when optimizing all these hyperparameters, simple logistic regression trained by stochastic gradient descent was performing as well as convolutional neural networks, which were the state of the art at the time of publication.
Hence, hyperparameters play a critical role in machine learning as they do not only determine the generalization capabilities of trained models, but they may actually determine what constitutes the state-of-the-art. Indeed, results reported in an empirical study might look very differently for another choice of hyperparameters and so would be the conclusions drawn. Unfortunately, it is not uncommon that published results do not report the specific hyperparameters that were used to run experiments, for instance, to demonstrate that a proposed method is superior to previously published ones, and are thus not reproducible, putting in question what actually is the state-of-the-art in machine learning today [haibe-kains:2020:transparency].
The subject of this monograph is hyperparameter optimization (HPO). The availability of algorithmic tools to automate this task and their adoption will help machine learning researchers dealing with hyperparameters in a principled way. More importantly, these tools will guarantee that new published advances are reproducible and thus unequivocally contribute to the field. Along with the experimental set-up, we argue that researchers would need to publish the HPO algorithm they used, the amount of resources it was allowed to consume [dodge:2019:show_your_work], and the hyperparameter search space that was considered [gundersen:2018:state_art]. Our objective is to provide a comprehensive overview of established techniques and recent advances in HPO. By doing so we hope we will help establish the practice of systematically using automatic HPO algorithms in research and industry.
@@ 1.1 What is hyperparameter optimization?
What distinguishes the HPO problem from conventional optimization problems arising for instance in convex or combinatorial optimization, is its nested nature: to evaluate the response function $f$ at a specific point $\lambda$, it is required to run the learning algorithm $A$. In practice, running $A$ can be time-consuming. For example, training ResNet-50 from scratch on CIFAR-10 on an Amazon Elastic Cloud Compute (EC2) g4dn.xlarge instance takes more than 2 hours. [^1] A ResNet is a deep residual neural network commonly used in computer vision. ResNet-50 has 50 convolutional layers with skip connections (one for every stack of three layers). We refer the interested reader to the original paper by he2016deep.
@@ 1.2 Why is hyperparameter optimization difficult?
In general, computing $f$ might not only be time-consuming, but also expensive in terms of amount of resources used or cost. Moreover, $f$ is required to satisfy minimal regularity constraints in practice. Its analytical form might not be known, it might have multiple local optima, it might exhibit wide flat regions, or it might be even non-smooth or non-continuous as a function of $\lambda$. We only require that $f$ is observable. Returning to the example of AlphaGo, this means that it does not matter whether the mathematical expression linking the hyperparameters of AlphaGo (i.e., how the algorithm is configured) and the probability of winning a game is known; what matters is that, for our specific choice of $\lambda$, we can observe if AlphaGo won or lost the game. Therefore, in general, one cannot apply conventional gradient-based optimization algorithms to solve HPO problems.
In addition, evaluations of $f$ can be further corrupted by noise, and the search space for $\lambda$ can itself be quite complex. Some hyperparameters could be continuous (e.g., learning rate), but others could be integral (e.g., number of layers in a deep neural network) or categorical to encode choices. Even more structure can emerge in the search space: conditional hyperparameters are variables that are only relevant depending on the value taken by others. For example, the number of units of the $l$-th layer of a neural network is only defined if the network has at least $l$ layers. If the domain to search over would be a small discrete set, one could carry out an exhaustive search by evaluating all the possible hyperparameter configurations in this set. Unfortunately, this is rarely the case in reality as the number of hyperparameter configurations to consider can be large or even uncountable.
All these challenges mean that, in practice, the focus of HPO is more often about finding “reasonably good” configurations as quickly as possible — which could still mean days of computation — rather than finding (global) optima. Hence, HPO often consists in determining a small set of good hyperparameters where to evaluate $f$.
@@ 1.4 Outline of the monograph
In this chapter, we gave a broad overview of the HPO problem, motivating its importance and its place in machine learning. In the next chapter, we will formalize many concept introduced here and further present several examples of hyperparameters and HPO problems. In Chapter 3, we will introduce elementary algorithms that search over a predefined or random grid of hyperparameters. While conceptually simple, random search strategies based on uniform sampling are attractive as they are embarrassingly parallel and exhibit an anytime behavior. However, they require the evaluation of a relatively large number hyperparameters to find a good solution. In Chapter 4, we will study model-based strategies, such as Bayesian optimization. These approaches learn a surrogate model of the response function $f(\lambda)$ based on the evaluations observed so far. Model-based approaches are sample-efficient, which means that they typically require a much smaller number of hyperparameters to be evaluated to find a satisfactory solution. However, they are inherently sequential, which can negatively impact the run time of the optimization algorithm. In Chapter 5, we will discuss how randomized approaches can be made resource-efficient by stopping unpromising evaluations early. We will also show that they can be married to model-based approaches in order to sample the hyperparameter search space in a non-uniform fashion. Chapter 6 covers population-based methods, which are genetic algorithms designed to solve global optimization problems. When applied to HPO, techniques in this family maintain a set of predictors whose hyperparameters can be updated dynamically, unlike the aforementioned approaches where hyperparameters are fixed for the entire training duration. In Chapter 7, we will discuss how gradient-based approaches can be applied to the optimization of continuous hyperparameters through hypergradients, which is more time-efficient than search-based algorithms. Finally, we close the monograph by discussing advanced use cases. In particular, we study how HPO is amenable to constrained and multi-objective settings. Indeed, HPO algorithms provide a powerful set of tools to meet competing requirements in applications (e.g., maximizing the predictive performance while mitigating the prediction bias) and satisfy practical constraints (e.g., maximal prediction latency).
@@ 1.5 Notation
As usual, we denote by $\mathbb{R}$ and $\mathbb{N}$ the set of real and natural numbers, respectively. For $K\in\mathbb{N}^{+}$, we indicate by $[K]=\{1,\dots,K\}$ the subset of the natural numbers from $1$ to $K$, included. We use calligraphic letters, like $\mathcal{D}$, to indicate sets, and script-style calligraphic, like $\mathscr{D}$ to indicate collections. We denote by $\mathcal{N}(\mu,\Sigma)$ the Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$.
@@
In this chapter we formalize hyperparameter optimization (HPO) along the concepts introduced in the previous chapter. We will make use of examples of HPO problems to fix ideas and showcase different application scenarios. We conclude the chapter with an overview of the practical desiderata for HPO that we will return to when discussing the various techniques covered in this monograph.
@@ 2.1 Learning algorithms with hyperparameters
Hyperparameters are variables that configure the behavior of a learning algorithm, shaping and modifying its inductive bias, i.e., the set of assumptions the algorithm relies on in order to generalize outside of the training data. We view a learning algorithm $A$ as a higher-order, possibly stochastic, function that maps data and hyperparameters to a model $h$, capable of accomplishing a given task:
$A:\mathscr{D}\times\Lambda\to\mathcal{H};\quad h_{\lambda}=A(\mathcal{D},\lambda),$ (2.1)
where:
  • $\mathscr{D}$ is the collection of all finite datasets that the algorithm can process;
@@
  • $\Lambda$ is the hyperparameter space, i.e., a subset of the Cartesian product of the domains of individual hyperparameters;
@@
  • $\mathcal{H}$ is the model space, typically indexed by a parameter vector $w$, that represents the class of models that $A$ may output.
@@
We call $\lambda\in\Lambda$ the hyperparameter vector or configuration. Note that in Eq. (2.1) we emphasize that the model returned by the algorithm depends on the hyperparameter vector $\lambda$ by writing $h_{\lambda}=A(\mathcal{D},\lambda)$. Likewise, in the following, we may omit that $A$ depends on the training data $\mathcal{D}$ when the context is clear. Indeed, typically, $\mathcal{D}$ is considered fixed in HPO.
2.1.1 Logistic regression for sentiment analysis

### v2

@@ Abstract
Hyperparameters are configuration variables controlling the behavior of machine learning algorithms. They are ubiquitous in machine learning and artificial intelligence and the choice of their values determines the effectiveness of systems based on these technologies. Manual hyperparameter search is often unsatisfactory and becomes infeasible when the number of hyperparameters is large. Automating the search is an important step towards advancing, streamlining, and systematizing machine learning, freeing researchers and practitioners alike from the burden of finding a good set of hyperparameters by trial and error. In this survey, we present a unified treatment of hyperparameter optimization, providing the reader with examples, insights into the state-of-the-art, and numerous links to further reading. We cover the main families of techniques to automate hyperparameter search, often referred to as hyperparameter optimization or tuning, including random and quasi-random search, bandit-, model-, population-, and gradient-based approaches. We further discuss extensions, including online, constrained, and multi-objective formulations, touch upon connections with other fields such as meta-learning and neural architecture search, and conclude with open questions and future research directions.
@@
Over the last decade, machine learning has established itself as a key driver of innovation in science and industry. The digitalization of human activities and interactions has led to the generation of data at an unprecedented scale. These data could be stored ever more cheaply thanks to advances in micro-electronics, as well as the advent of cloud computing. Progress in hardware and low-power chip design also led to an exponential increase of computing capabilities of respectively cloud infrastructure, notably graphical processing units (GPUs), and devices on the edge, such as mobile phones. These two trends, cheaper and better hardware for storage and compute, combined with strong recent advances in algorithms and architectures for deep learning, lay the foundation for the successful application of statistical machine learning in practice and its adoption in artificial intelligence. It enabled learning algorithms to sift through vast amounts of data to discover and extract patterns that can then be used to make data-driven decisions without or with limited human intervention.
@@
One notable illustration of the recent advances made in machine learning is AlphaGo [silver:2017:mastering_go], a computer program developed by machine learning experts at the London-based company DeepMind. AlphaGo won against one of the world champions of the game of Go, Lee Sedol, in 2016. This event generated considerable press coverage (and was even made into a movie) as it was believed at the time that no computer program could win a game of Go against a human before several decades to come. What is less known is that the success of AlphaGo critically depended on how a set of configuration parameters, called hyperparameters, were adjusted automatically by another computer program [chen:2018:bayesian_optimization_alphago]. This computer program relied on Bayesian optimization, a technique that sequentially predicts and then evaluates the performance of algorithms such as AlphaGo when it uses specific hyperparameters. Until a certain cost or time budget is exhausted, Bayesian optimization repeatedly trades off exploration and exploitation to determine what is the next most interesting set of hyperparameters to evaluate based on the system performances observed so far.
@@
In practice, the generalization capabilities of any machine learning algorithm also depend on hyperparameters. Whereas conventional (shallow) algorithms such as support vector machines [shawe-taylor:2004:kernel_methods_pattern_analysis] need hand-engineered features to transform the raw data in a suitable format, deep neural networks learn in an “end-to-end” fashion by ingesting the raw data directly [lecun2015nature]. However, in both cases, the composition of the model (e.g., the choice of feature extractors or the neural network architecture) and the empirical loss minimizer depend on configuration parameters that are not adjusted by the algorithm itself. Ultimately, their choice contributes to determine how effectively the algorithm can learn a specific natural language understanding, computer vision, or speech recognition task.
1.1 A simple illustrative example
@@
To illustrate the practical importance of carefully selecting the hyperparameters, we consider the problem of sentiment analysis. yogotama2015-emnlp15 studied the effect of hyperparameters in this context. More specifically, the authors framed the problem as a standard binary classification problem as is commonly done in the literature, where the task of the classifier is to predict whether the text expresses a negative or a positive sentiment. They compared simple logistic regression trained with stochastic gradient descent to convolutional neural networks, which achieved state-of-the-art results at the time of the publication. Table 1.1(a) shows the hyperparameters that were searched over. These included the choice of text features used (e.g., removal of stop words or not), the type of regularization (e.g., $\ell_{1}$ or $\ell_{2}$) and optimization algorithm parameters (e.g., convergence tolerance). The experimental results they reported on the Amazon electronics dataset are reproduced in Table 1.1(b). Interestingly, they showed that tuned logistic regression using a bag of words representation of the text performed on par with a convolutional neural network and only slightly worse than a sequential convolutional neural network. Hence, by carefully optimizing hyperparameters, the authors could achieve a result close to the state of the art with a much simpler model. We will come back to this example in more details in the next chapter.
@@
Figure 1.1: Results on sentiment analysis reported in [yogotama2015-emnlp15]. Hyperparameters that were optimized and associated domains are listed on the left table. The feature extraction depends on the range of $n$-grams considered through $n_{min}$ and $n_{max}$. Stop words could be removed or not and words weighted according different weighting schemes. Two choices of regularization were considered and the amount of regularization was optimized. The convergence tolerance was also optimized. The binary classifiers compared are either SVM (support vector machine), LR (logistic regression), NN (multi-layer perceptron) or CNN (convolutional neural network). The authors found that when optimizing all these hyperparameters, simple logistic regression trained by stochastic gradient descent was performing as well as convolutional neural networks, which were the state of the art at the time of publication.
Hyperparameters play a critical role in machine learning as they do not only determine the generalization capabilities of trained models, but they may actually determine what constitutes the state-of-the-art. Indeed, results reported in an empirical study might look very differently for another choice of hyperparameters and so would be the conclusions drawn. Unfortunately, it is not uncommon for published results to omit specific hyperparameters used in the experiments, which may hinder reproducibility. Moreover, when experiments support the claim that a proposed method surpasses previous ones, the instability of results with respect to hyperparameters may cast doubts on these claims and raise questions about the current state-of-the-art [haibe-kains:2020:transparency].
The subject of this monograph is hyperparameter optimization (HPO). The availability of algorithmic tools to automate this task and their adoption will help machine learning researchers dealing with hyperparameters in a principled way. Importantly, these tools will also significantly help towards ensuring that new published advances are reproducible and thus unequivocally contribute to the field. Along with the experimental set-up, we argue that researchers would need to publish the HPO algorithm they used, the amount of resources it was allowed to consume [dodge:2019:show_your_work], and the hyperparameter search space that was considered [gundersen:2018:state_art]. Our objective is to provide a comprehensive overview of established techniques and recent advances in HPO. By doing so we hope we will help establish the practice of systematically using automatic HPO algorithms in research and industry, besides offering an introductory guide to understanding the fundamentals of this field.
@@ 1.2 What is hyperparameter optimization?
What distinguishes the HPO problem from conventional optimization problems, such as those arising in convex or combinatorial optimization, is its nested nature: to evaluate the response function $f$ at a specific point $\lambda$, it is required to run the learning algorithm $A$, which in turn has to solve another optimization problem to fit a model to the available data. In practice, running $A$ can be time-consuming. For example, training ResNet-50 from scratch on CIFAR-10 on an Amazon Elastic Cloud Compute (EC2) g4dn.xlarge instance takes more than 2 hours. [^1] A ResNet is a deep residual neural network commonly used in computer vision. ResNet-50 has 50 convolutional layers with skip connections (one for every stack of three layers). We refer the interested reader to the original paper by he2016deep.
@@ 1.3 Why is hyperparameter optimization difficult?
In general, computing $f$ might not only be time-consuming, but also expensive in terms of amount of resources used or cost. Its analytical form is typically either unknown or impractical to handle, it might have multiple local optima, it might exhibit wide flat regions, or it might be even non-smooth or non-continuous as a function of $\lambda$. As we will see in the following, for some classes of algorithms, we only require that $f$ is observable. Returning to the example of AlphaGo, this means that it does not matter whether the mathematical expression linking the hyperparameters of AlphaGo (i.e., how the algorithm is configured) and the probability of winning a game is known; what matters is that, for our specific choice of $\lambda$, we can observe if AlphaGo won or lost the game. Therefore, in general, one cannot apply conventional gradient-based optimization algorithms to solve HPO problems (but we will see that with additional assumption on $f$ this is, in fact, possible).
In addition, evaluations of $f$ can be stochastic, and the search space for $\lambda$ can itself be quite complex. Some hyperparameters could be continuous (e.g., learning rate), but others could be integral (e.g., number of layers in a deep neural network) or categorical to encode choices. Even more structure can emerge in the search space: conditional hyperparameters are variables that are only relevant depending on the value taken by others. For example, the number of units of the $j$-th layer of a neural network is only defined if the network has at least $j$ layers. If the domain to search over would be a small discrete set, one could carry out an exhaustive search by evaluating all the possible hyperparameter configurations in this set. Unfortunately, this is rarely the case in reality as the number of hyperparameter configurations to consider can be large or even uncountable.
All these challenges mean that, in practice, the focus of HPO is more often about finding “reasonably good” configurations as quickly as possible — which could still mean days of computation — rather than finding (global) optima. Hence, HPO often consists in determining a small set of good hyperparameters on which to evaluate $f$.
@@ 1.5 Outline of the monograph
In this chapter, we gave a broad overview of the HPO problem, motivating its importance and its place in machine learning. In the next chapter, we will formalize many concept introduced here and further present several examples of hyperparameters and HPO problems. In Chapter 3, we will introduce elementary algorithms that search over a predefined or random grid of hyperparameters. While conceptually simple, random search strategies based on uniform sampling are attractive as they are embarrassingly parallel and exhibit an anytime behavior. However, they require the evaluation of a relatively large number hyperparameters to find a good solution. In Chapter 4, we will study model-based strategies, such as Bayesian optimization. These approaches learn a surrogate model of the response function $f(\lambda)$ based on the evaluations observed so far. Model-based approaches are sample-efficient, which means that they typically require a much smaller number of hyperparameters (compared to grid or random search) to be evaluated to find a satisfactory solution. However, they are inherently sequential, which can negatively impact the run time of the optimization algorithm. In Chapter 5, we will discuss how randomized approaches can be made resource-efficient by stopping unpromising evaluations early. We will also show that they can be married to model-based approaches in order to sample the hyperparameter search space in a non-uniform fashion. Chapter 6 covers population-based methods, stochastic algorithms for global optimization that draw inspirations from natural phenomena. When applied to HPO, some techniques in this family such as population-based training maintain a set of predictors whose hyperparameters are updated dynamically, unlike the aforementioned approaches where hyperparameters are fixed for the entire training duration. In Chapter 7, we will discuss how gradient-based approaches can be applied to the optimization of continuous hyperparameters through hypergradients, which is more time-efficient than search-based algorithms. Finally, we close the monograph by discussing advanced use cases. In particular, we study how to adapt HPO to constrained and multi-objective settings. Indeed, HPO algorithms provide a powerful set of tools to meet competing requirements in applications (e.g., maximizing the predictive performance while mitigating the prediction bias) and satisfy practical constraints (e.g., maximal prediction latency). Furthermore, we also discuss connections with with neural architecture search and meta-learning, and touch upon transfer learning of hyperparameters for foundation models. The two closing chapters are dedicated to an overview of HPO software frameworks (Chapter 9) and emerging topics and open questions in the field (Chapter 10).
1.6 Related surveys
The topic of this monograph has been surveyed in some other papers, briefly in [hutter_beyond_2015] and more extensively in  [bischl:2023:hyperparameter_optimization_foundations, yang:2020:hyperparameter_optimization_machine]. Arlot and Celisse arlot:2010:survey_crossvalidation_procedures surveyed the model selection problem from a complementary statistical perspective. Specific aspects have been reviewed in other review papers, in particular we would like to mention Bayesian optimization [Frazier:08, shahriari:2016:taking_human_out, wang:2023:recent_advances_bayesian], neural architecture search [ren:2021:comprehensive_survey_neurala, heuillet:2024:efficient_automation_neural], evalutionary algorithms [li:2023:survey_evolutionary_deep], and meta-learning [hospedales2020meta] HPO is a core topic in automated machine learning (AutoML). The subject is covered extensively in the book by Hutter et al. hutter2019automated and more recently in [he:2021:automl]
@@ 1.7 Notation
As usual, we denote by $\mathbb{R}$ and $\mathbb{N}$ the set of real and natural numbers, respectively. For $K\in\mathbb{N}^{+}$, we indicate by $[K]=\{1,\dots,K\}$ the subset of the natural numbers from $1$ to $K$, included. We use calligraphic letters, like $\mathcal{D}$, to indicate sets, and script-style calligraphic, like $\mathscr{D}$ to indicate collections. We reserve the Greek letter $\lambda$ to indicate hyperparameters and $f$ for the response function. We use $\mathbb{P}$ to indicate probability functions. Whenever confusion may arise, we further add a subscript like $\mathbb{P}_{x\sim\nu}$ to specify that the random variable $x$ evaluated by $\mathbb{P}$ follow the distribution $\nu$. We use $\mathbb{E}$ with the same convention to indicate the expected value. We mostly reserve the letter $p$ to indicate density or probability mass functions. We denote by $\mathcal{N}(\mu,\Sigma)$ the Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$.
@@
Here, we formalize hyperparameter optimization (HPO) along the concepts introduced in the previous chapter. We discuss the nested structure that characterizes an HPO problem and the archetypal settings encountered in machine learning, contrasting this to more general algorithm configuration problems. To fix ideas, we revisit in greater detail the example of the introduction and showcase different application scenarios from the literature. We conclude the chapter with an overview of practical desiderata for HPO algorithms that will guide our discussions in the remainder of the monograph.
@@ 2.1 Learning algorithms with hyperparameters
Hyperparameters are variables that configure the behavior of a learning algorithm, shaping and modifying its inductive bias, i.e., the set of assumptions the algorithm relies on in order to generalize outside of the training data. Formally, let:
  • $\mathscr{D}$ be a the collection of all finite datasets that the algorithm can process;
@@
  • $\Lambda$ be an hyperparameter space, often though of as the Cartesian product of individual hyperparameters spaces;
@@
  • $\mathcal{H}$ be a model space, typically indexed by a parameter vector $w$, that represents the class of models that the algorithm may output.
We view a learning algorithm $A$ as a higher-order, possibly stochastic, function that maps data and hyperparameters to a model $h_{\lambda}$, capable of accomplishing a given task:
$A:\mathscr{D}\times\Lambda\to\mathcal{H};\quad h_{\lambda}=A(\mathcal{D},\lambda).$ (2.1)
@@
We invariably call a $\lambda\in\Lambda$ a hyperparameter vector or a hyperparameter configuration. By writing $h_{\lambda}=A(\mathcal{D},\lambda)$ in Eq. (2.1) we emphasize the dependence of the model returned by the algorithm on the hyperparameter vector $\lambda$, but omit the dependence on the training data $\mathcal{D}$. This choice is consistent with the fact that often one considers the dataset fixed and “given” when setting up an HPO problem.
2.1.1 Logistic regression for sentiment analysis: an HPO perspective